An alternative method for preparing pure iron from Fe2O3 is by reaction with carbon monoxide.
Fe2O3 + CO Fe + CO2 (Unbalanced)

(a) Balance the equation. (Use the lowest possible coefficients.)

Fe2O3 +
CO
Fe +
CO2
(b) How many grams of CO are needed to react with 3.15 g of Fe2O3?

g
(c) How many grams of CO are needed to react with 1.75 mol of Fe2O3?

g
1
Expert's answer
2018-01-24T15:10:08-0500
a)Fe2O3 + 3CO = 2Fe + 3CO2

b) Mr(Fe2O3) = 160g/mol
Mr(CO) = 28 g/mol
m(CO) = 3.15*84/160 = 1.65(g)

c) m(CO) = 84*1.75 = 147(g)